jman12354:

One side of a square is 3x + 2. Which expression represents the perimeter?

Tranquility:

Hello! Welcome to QuestionCove! Please post your complete question along with answer choices to receive better assistance.

Tranquility:

Regardless, we can still answer your question.. The perimeter of a square is 4 * side we know that the side length is (3x+2) The perimeter is the sum of all the sides of the object, which is a square in this side. Since a square has 4 sides and all sides are the same length, we just multiply the side length by 4 to find the perimeter. What do you get? Also, remember the distributive property a(b + c) = ab + ac a(b - c) = ab - ac